# Enabling Weak Client Participation via On-Device Knowledge Distillation in Heterogeneous Federated Learning

### v1

@@ 1 Introduction
Federated Learning (FL) [1] is a practical learning method to utilize the data collected on-the-fly at the edge. Recently, with the rapid advancement of edge device system efficiency, various large-scale on-device learning scenarios are being considered, such as multi-task transfer learning [2], sharpness-aware minimization [3], and even Large Language Model (LLM) fine-tuning [4]. These on-device learning studies commonly assume that the devices are homogeneous and thus all the devices can equally contribute to the model training. However, FL environments often consist of heterogeneous edge devices such as mobile phones or IoT devices. Therefore, to exploit the ever-increasing big data collected at the edge, it is imperative to develop a new FL method that can fully utilize heterogeneous system resources. Several recent studies tackle the system heterogeneity issue in FL by employing partial model training scheme [5, 6, 7, 8, 9, 10]. However, they enforce clients with weak resources to train only a subset of the target model’s layers or channels, significantly harming the freedom of model design. More recent works utilize Knowledge Distillation (KD) [11]. They commonly use a logit ensemble method instead of directly aggregating the local models to reduce the KD’s communication cost in FL environments [12, 13]. In our empirical study, however, we find that the logit ensemble fails to provide sufficiently accurate soft targets when the data distribution is highly non-IID. Although a large amount of public data is available, KD with such inaccurate soft targets negatively impacts the efficiency of knowledge transfer. Some others utilize unlabeled extra data to train the target model on the server side [14, 15]. This approach assumes that unlabeled data generated on edge devices is aggregated on the server and utilized for centralized training, raising privacy concerns.
@@
In this paper, we propose a novel heterogeneous FL method that utilizes unlabeled data at the edge to overcome the limitations of logit ensemble methods and server-side model training approaches. Specifically, we introduce a KD-based FL method that jointly utilizes additional system resources on strong devices and unlabeled local data by combining auxiliary model training and on-device KD. The core principle behind our method is that a small auxiliary model first learns global knowledge through standard model aggregation and then helps the training of the large target model by utilizing system resources and unlabeled local data on strong devices. When transferring global knowledge from the auxiliary model to the target model, only a subset of clients with sufficiently strong system resources participate in the KD process. This heterogeneous approach allows all the clients effectively join the training exploiting their own resources and local data.
@@
Compared to logit ensemble-based KD, our method has two strong benefits as follows. First, our approach ensures accurate soft targets for the KD process. By aggregating the small auxiliary models across the clients, the local knowledges are directly merged into a single point, and it results in making the quality of the soft targets remarkably improved compared to the logit ensemble. Second, our approach does not require the server to have sufficiently strong system capability nor the unlabeled data to be centralized on the server in advance. By offloading the KD workload to the strong devices, the model can learn from the local data regardless of whether it is labeled or not. Moreover, as more clients join the training, our approach allows the model to learn from more unlabeled data, likely improving the model accuracy.
@@
  • We explain and empirically prove the fundamental limitations of the logit ensemble-based Knowledge Distillation especially when the data is strongly non-IID.
@@ 2.2 Knowledge Distillation in Federated Learning
Logit Ensemble, Server-Side Distillation, and Heterogeneous Models – FedGKT [28] allows the server to have its own classifier layers and train them using KD. This approach achieves promising model accuracy, however, it does not take the public data into account. FedGEMS [29] also runs KD to train a large model at the server-side. Likely to FedMD, this method also assumes the public data to be labeled. More recently, Fed-ET [15] has been proposed, which utilizes model transfer as well as weighted ensemble at the server-side. However, this approach also performs KD on the server side, and the server-side model cannot directly learn from the clients’ private data.
Table 1: The configurations of SOTA FL methods.
@@
FedGEMS $>$ client logits Yes / Yes both
FedGKT $>$ client model+logits No / N/A both
Fed-ET $>$ client model+logits Yes / No server
@@ 3 Method
We begin with an ablation study to demonstrate how logit ensemble can harm KD performance in general FL environments. Based on this analysis, we develop our novel on-device KD-based FL method, which maximizes KD performance by fully utilizing the system resources of heterogeneous devices.
@@ 3.1 Limitations of Logit Ensemble in Federated Learning Environments
We compare the accuracy of logit ensemble and that of the conventional model aggregation. We use CIFAR-10 and Google Speech Command, computer vision and audio recognition benchmarks. First, the dataset is partitioned into two distinct subsets: one for training and the other for evaluating model accuracy. We further divide the training data into 100 small subsets using a label-based Dirichlet distribution ($\alpha=0.1$) and assign one small subset to each client. Next, we conduct logit ensemble-based FL and model aggregation-based FL, and compare the accuracies they achieve. Note that the logit ensemble uses the full model, while the model aggregation uses a smaller model, with each layer containing $25\%$ of the channels. This setting penalizes the model aggregation scheme in terms of the model’s learning capability, and we expect to see the adverse impact of the logit ensemble more clearly under this setting.
Figure 1: Non-IID distributions of CIFAR-10 (left-top) and Google Speech (left-bottom). Label-based Dirichlet distributions ($\alpha=0.1$) are used to get 100 local distributions. Due to the limited space, we only show 20 of them. The accuracy of small model trained via FL provides remarkably more accurate logits than the logit ensemble.
@@ 3.2 On-Device Federated Transfer Learning
We focus on federated optimization, which trains a target model $\textbf{w}_{l}$. Considering the natural data collection at the edge, we assume that each edge device has a certain amount of unlabeled data. This is a valid problem formulation because it is more reasonable for new data to be generated at the edge rather than the parameter server. Our method employs KD with these unlabeled local data to improve the target model accuracy. We take advantage of an auxiliary model $\textbf{w}_{s}$, which fits within the memory capacity of all individual devices. Given the two models, our method repeats the following two steps until a user-defined termination condition is satisfied (e.g., target accuracy or time budget). Figure 2 provides a schematic illustration of these two steps.
Step 1: Auxiliary Model Training Round – First, a random subset of clients is activated in each round. The activated devices receive $\textbf{w}_{s}$ from the server and run $\tau$ local updates using their own labeled data. Then, the locally trained models, $\textbf{w}_{s}^{i}$ for $i\in A$, are aggregated at the server, where $A$ is the set of active clients. This global FL round with the small auxiliary model enables the collection of local knowledge from all clients, regardless of their system capacity. The loss function in step 1 is as follows.
@@
Step 2: Target Model Training Round – Once the auxiliary models are aggregated at the server, the strong clients train the target model $\textbf{w}_{l}$ using the auxiliary model. First, the strong clients receive both models, $\textbf{w}_{s}$ and $\textbf{w}_{l}$ from the server. Then, they locally update the target model $\textbf{w}_{l}$ for $\tau$ steps using their own labeled data. After the local updates, they run additional $\tau$ local KD steps using the unlabeled local data. The auxiliary model participates in the KD process as a teacher. Finally, the local target models $\textbf{w}_{l}^{j},\forall j\in S$ are aggregated at the server, where $S$ is the set of strong clients.
@@
Figure 2: The schematic illustration of the proposed heterogeneous FL method. For simplicity, the server is omitted in this schematic. An auxiliary (small) model is trained by all the devices using their private data. Then, the global knowledge is transferred to the target (large) model through on-device KD. This distributed knowledge transfer approach enables to not only fully utilize heterogeneous edge devices’ system resources but also effectively use the unlabeled data for training.
@@ 3.3 Discussion on Design Choices of On-Device Knowledge Distillation
In addition, the logit ensemble allows local models to remain personalized, as model parameters are not directly averaged across clients. If data is highly non-IID – for example, if each client has local data limited to one or two classes – these personalized models are likely to suffer from poor generalization performance. Our ablation study shows in Figure 1. Thus, we can conclude that aggregating local logits alone is insufficient to provide accurate soft targets for KD.
@@
Communication Cost Analysis – We compare the communication cost of our method with that of representative logit ensemble-based KD methods. Table 2 shows the results of the comparative study using CIFAR-10 (ResNet20). FedDF employs server-side KD, whereas DS-FL adopts client-side KD. To ensure a fair comparison, we align the total number of rounds, regardless of whether they involve training or KD.
For the homogeneous FL methods (FedAvg and FedDF), we consider two cases: one case is when the small model $w_{s}$ (with $25\%$ of the channels at every layer) is trained on all clients, and the other case is when the large model $w_{l}$ is trained only on a subset of clients with sufficiently strong system resources. First, FedAvg (small) has a lower communication cost than FedAvg (large), but suffers from significantly lower accuracy. Interestingly, the accuracy of FedDF is lower than that of FedAvg, regardless of the model size. This result indicates that the soft targets generated by the logit ensemble are of insufficient quality to make meaningful progress in KD. FedDF can leverage KD without incurring extra communication cost. However, it runs KD on the server-side with inaccurate soft targets, which could ultimately harm model accuracy. DS-FL, the client-side KD-based FL method, adopts the logit ensemble and entirely avoids the model aggregation cost. Thus, the communication cost is only the number of classes $C$, but the model accuracy is significantly reduced.
@@
[…]
FedAvg(small) $|w_{s}|$ 19K $42.22\pm 0.1\%$
@@ 3.4 Discussion on Impact of On-Device Knowledge Distillation on Generalization Performance
Here, we analyze the benefit of our design choices in terms of the target model’s generalization bound. Our analysis builds upon that of [15], extending it to account for the combination of labeled and unlabeled local data.
Problem Definition – We first define hypothesis $h:\mathcal{X}\rightarrow\mathcal{Y}$, with input space $\mathbf{x}\in\mathcal{X}$ and label space $y\in\mathcal{Y}$, and the hypothesis space $h\in\mathcal{H}$. The general loss function $l(h(\mathbf{x}),y)$ quantifies the classification error of the given hypothesis. The true global data distribution is defined as $\mathcal{D}$. Then, the true local data distribution from each client’s perspective is defined as $\mathcal{D}_{i},i\in[n]$. The data distribution of the limited number of local observations is defined as $\mathcal{\hat{D}}_{i}$ We also define the expected loss over an arbitrary data distribution $\mathcal{D}^{\prime}$ as $\mathcal{L}_{\mathcal{D^{\prime}}}(h)=\mathbb{E}_{\mathbf{x},y\sim\mathcal{D}^{\prime}}\left[l(h(\mathbf{x}),y)\right],\forall h\in\mathcal{H}$. Finally, we assume that $\mathcal{L}(h)$ is convex. In these settings, the generalization bound of the target model trained by our proposed method satisfies the following condition.
To make our analysis self-contained, we state the two lemmas proposed in the previous work, which will be used to analyze the generalization of our proposed method.
Lemma 1 (Domain Adaptation) [30].
With two true distributions $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, $\forall p\in(0,1)$ and hypothesis $\forall h\in\mathcal{H}$, with probability at least $1-p$ over the choice of samples, it satisfies
$\mathcal{L}_{\mathcal{D}_{1}}(h)\leq\mathcal{L}_{\mathcal{D}_{2}}(h)+\frac{1}{2}d(\mathcal{D}_{1},\mathcal{D}_{2})+v,$ (3)
where $d(\mathcal{D}_{1},\mathcal{D}_{2})$ quantifies the distribution discrepancy between the two data distributions, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, $v_{i}$ is the minimal loss such that $v_{i}=\textrm{inf}_{h}\mathcal{L}_{\mathcal{D}_{i}}(h)+\mathcal{L}_{\mathcal{D}}(h)$.
@@
The first lemma provides a generalization bound for one data distribution based on the maximum discrepancy between the two given distributions. The second lemma refines this bound by considering the limited training samples, leveraging Hoeffding’s inequality. Our proposed method trains the target model $\mathbf{x}_{l}$ using only $s<n$ strong clients. Accordingly, we analyze the generalization bound with respect to the labeled data from the strong clients using the two lemmas above. Subsequently, we revise this bound to incorporate on-device KD with unlabeled local data. See Appendix for proof.
Proposition 1 Consider $n$ clients in total and $s$ of them are strong clients. Each client has its own data distribution $\mathcal{D}_{i},\forall i\in[n]$. We define a combination of the labeled local data and the unlabeled local data as $\tilde{\mathcal{D}}_{i}$. For an arbitrary data distribution $\mathcal{D}^{\prime}$, $h_{\mathcal{D}^{\prime}}$ is the model best-trained on $\mathcal{D}^{\prime}$. Then, with probability of $1-p$, our proposed on-device KD-based FL guarantees the following generalization bound for the target model.
$\mathcal{L}_{\mathcal{D}}\left(\frac{1}{s}\sum_{i=1}^{s}h_{\mathcal{\tilde{D}}_{i}}\right)$ $\leq\frac{1}{s}\sum_{i=1}^{s}\left(\mathcal{L}_{\tilde{\mathcal{D}}_{i}}\left(h_{\mathcal{\tilde{D}}_{i}}\right)+\sqrt{\frac{\textrm{log}2/p}{2|\tilde{\mathcal{D}_{i}}|}}\right)$
@@
Remark 1. This result shows the key advantage of our method: The on-device KD reduces the second term on the right-hand side in (6), which means better generalization. Remember that $|\mathcal{\hat{D}}_{i}|\leq|\mathcal{\tilde{D}}_{i}|$, where $\hat{\mathcal{D}}_{i}$ is the labeled data and $\tilde{\mathcal{D}}_{i}$ is the combination of labeled and unlabeled data. Thus, we can expect that the proposed method achieves a lower empirical loss by enlarging the data distribution.
Remark 2. As each client collects more unlabeled data, the size of $\mathcal{\tilde{D}}_{i}$ grows, and the bound becomes even smaller. This is a crucial benefit in realistic FL environments, as new unlabeled data is most likely collected at the edge rather than the server. So, our method can be considered as a more scalable FL method especially when the strong clients can collect data on-the-fly. Note that in the server-side KD, the $\mathcal{\hat{D}}_{i}$ can only be the unlabeled data which is still smaller than $\mathcal{\tilde{D}}_{i}$.

### v2

@@ 1 Introduction
Federated Learning (FL) [26] is a practical learning method to utilize the data collected on-the-fly at the edge. With recent advances in the system efficiency of edge devices, various large-scale on-device learning scenarios are being considered, such as multi-task transfer learning [6], sharpness-aware minimization [11], and even Large Language Model (LLM) fine-tuning [35]. These studies generally assume homogeneous devices, where all participants can contribute equally to model training. However, FL environments often consist of heterogeneous edge devices, such as mobile phones and IoT devices, which continuously gather new, unlabeled local data [29, 5] that could be used for training. Therefore, to exploit the ever-increasing big data collected at the edge, it is imperative to develop a new FL method that can fully utilize heterogeneous system resources.
Several recent studies tackle the system heterogeneity issue in FL by employing partial model training scheme [9, 15, 20, 19, 21, 34]. However, they enforce clients with weak resources to train only a subset of the target model’s layers or channels, significantly harming the freedom of model design. More recent works utilize Knowledge Distillation (KD) [14]. They commonly use a logit ensemble method instead of directly aggregating the local models to reduce the KD’s communication cost in FL environments [17, 16]. In our empirical study, however, we find that the logit ensemble fails to provide sufficiently accurate soft targets when the data distribution is highly non-IID. Although a large amount of public data is available, KD with such inaccurate soft targets negatively impacts the efficiency of knowledge transfer. Some others utilize unlabeled extra data to train the target model on the server side [24, 8]. This approach assumes that unlabeled data generated on edge devices is aggregated on the server and utilized for centralized training, raising privacy concerns.
@@
In this paper, we propose a novel heterogeneous FL method that utilizes unlabeled data at the edge to overcome the limitations of logit ensemble methods and server-side model training approaches. Specifically, we introduce a KD-based FL method that jointly utilizes additional system resources on strong devices and unlabeled local data by combining auxiliary model training and on-device KD. The core principle behind our method is that a small auxiliary model first learns global knowledge through standard model aggregation and then helps the training of the large target model by utilizing system resources and unlabeled local data on strong devices. When transferring global knowledge from the auxiliary model to the target model, only a subset of clients with sufficiently strong system resources participate in the KD process. This heterogeneous approach allows all clients to effectively join the training, exploiting their own resources and local data.
@@
Compared to logit ensemble-based KD, our method has two strong benefits as follows. First, our approach ensures accurate soft targets for the KD process. By aggregating the small auxiliary models across clients, the local knowledge is merged into a single point, thereby significantly improving the quality of the soft targets compared to the logit ensemble. Second, our approach does not require the server to have sufficiently strong system capability nor the unlabeled data to be centralized on the server in advance. By offloading the KD workload to the strong devices, the model can learn from the local data regardless of whether it is labeled or not. Moreover, as more clients join the training, our approach allows the model to learn from more unlabeled data, likely improving the model accuracy.
@@
  • We explain and empirically prove the fundamental limitations of the logit ensemble-based Knowledge Distillation especially when data is strongly non-IID.
@@ 2.2 Knowledge Distillation in Federated Learning
Logit Ensemble, Server-Side Distillation, and Heterogeneous Models – FedGKT [12] allows the server to have its own classifier layers and train them using KD. This approach achieves promising model accuracy, however, it does not take into account locally generated unlabeled data. FedGEMS [7] also runs KD to train a large model at the server-side. Like FedMD, this method also assumes the public data to be labeled. More recently, Fed-ET [8] has been proposed, which utilizes model transfer as well as weighted ensemble at the server-side. However, this approach also performs KD on the server side, and the server-side model cannot directly learn from the clients’ private data.
@@
FedGEMS >> client logits Yes / Yes both
FedGKT >> client model+logits No / N/A both
Fed-ET >> client model+logits Yes / No server
@@ 3 Method
We begin with an ablation study to demonstrate how logit ensemble can harm KD performance in general FL environments. Based on this analysis, we develop our novel on-device KD-based FL method, which maximizes KD performance by fully utilizing the system resources of heterogeneous devices. Finally, we analyze the expected generalization bound for the proposed method and discuss how unlabeled local data impacts its performance.
@@ 3.1 Limitations of Logit Ensemble in Federated Learning Environments
We compare the accuracy of logit ensemble and that of the conventional model aggregation. We use CIFAR-10 and Google Speech Command, which are computer vision and audio recognition benchmarks. First, the dataset is partitioned into two distinct subsets: one for training and the other for evaluating model accuracy. We further divide the training data into 100 small subsets using a label-based Dirichlet distribution ($\alpha=0.1$) and assign one small subset to each client. Next, we conduct logit ensemble-based FL and model aggregation-based FL, and compare the accuracies they achieve. Note that the logit ensemble uses the full model, while the model aggregation uses a smaller model, with each layer containing $25\%$ of the channels. This setting penalizes the model aggregation scheme in terms of the model’s learning capability, and we expect to see the adverse impact of the logit ensemble more clearly under this setting.
@@
Figure 1: Non-IID distributions of CIFAR-10 (left-top) and Google Speech (left-bottom). Label-based Dirichlet distributions ($\alpha=0.1$) are used to get 100 local distributions. Due to the limited space, we only show 20 of them. The accuracy of a small model trained via FL provides remarkably more accurate logits than the logit ensemble.
Figure 2: The schematic illustration of the proposed heterogeneous FL method. For simplicity, the server is omitted in this schematic. An auxiliary (small) model is trained by all the devices using their private data. Then, the global knowledge is transferred to the target (large) model through on-device KD. This distributed knowledge transfer approach enables to not only fully utilize heterogeneous edge devices’ system resources but also effectively use the unlabeled private data for training.
@@ 3.2 On-Device Federated Transfer Learning
We first define two types of models: $\textbf{w}_{l}$ and $\textbf{w}_{s}$.
  • $\textbf{w}_{l}$: The large target model.
  • $\textbf{w}_{s}$: A smaller auxiliary model.
We also define two relative categories of clients: strong and weak clients. A client is considered strong if it has sufficient system resources to train the large target model $w_{l}$. On the other hand, a client is considered weak if it cannot independently train the large target model $w_{l}$. This limitation may be due to various system constraints, such as insufficient memory or too slow computational resources. We do not specify the exact reason in order to keep the definition of this type general.
Now, we focus on federated optimization, which trains the target model $\textbf{w}_{l}$. Considering the natural data collection at the edge, we assume that each edge device has a certain amount of unlabeled prviate data. This is a valid problem formulation because it is more reasonable for new data to be generated at the edge rather than the parameter server. Our method employs KD with these unlabeled local data to improve the target model accuracy. We take advantage of an auxiliary model $\textbf{w}_{s}$, which fits within the memory capacity of all individual devices.
Given the two models, our method repeats the following two steps until a user-defined termination condition is satisfied (e.g., target accuracy or time budget). Figure 2 provides a schematic illustration of these two steps.
Step 1: Auxiliary Model Training Round – First, a random subset of clients is activated in each round. The activated devices receive $\textbf{w}_{s}$ from the server and run $\tau$ local updates using their own labeled data. The resulting locally trained models, $\textbf{w}_{s}^{i}$ for $i\in A$, are aggregated at the server using simple uniform averaging, where $A$ is the set of active clients. This global FL round with the small auxiliary model enables the collection of local knowledge from all clients, regardless of their system capacity. The loss function in step 1 is as follows.
@@
Step 2: Target Model Training Round – Once the auxiliary models are aggregated at the server, the strong clients train the target model $\textbf{w}_{l}$ using the auxiliary model. First, the strong clients receive both models, $\textbf{w}_{s}$ and $\textbf{w}_{l}$ from the server. Then, they locally update the target model $\textbf{w}_{l}$ for $\tau$ steps using their own labeled data. After the local updates, they run additional $\tau$ local KD steps using the unlabeled private data. The auxiliary model participates in the KD process as a teacher. Finally, the local target models $\textbf{w}_{l}^{j},\forall j\in S$ are aggregated at the server, where $S$ is the set of strong clients.
@@ 3.3 Discussion on Design Choices of On-Device Knowledge Distillation
In addition, the logit ensemble allows local models to remain personalized, as model parameters are not directly averaged across clients. If data is highly non-IID – for example, if each client has local data limited to one or two classes – these personalized models are likely to suffer from poor generalization performance. Our ablation study results are shown in Figure 1. Thus, we can conclude that aggregating local logits alone is insufficient to provide accurate soft targets for KD.
@@
Communication Cost Analysis – We compare the communication cost of our method with that of representative logit ensemble-based KD methods. Table 2 presents the results of the comparative study conducted on CIFAR-10 (ResNet20). FedDF employs server-side KD, whereas DS-FL adopts client-side KD. To ensure a fair comparison, we align the total number of rounds, regardless of whether they involve training or KD.
For the homogeneous FL methods (FedAvg and FedDF), we consider two cases: one case involves training the small model $w_{s}$ (with $25\%$ of the channels at every layer) on all clients, while the other case trains the large model $w_{l}$ only on a subset of clients with sufficiently strong system resources. First, FedAvg (small) has a lower communication cost than FedAvg (large), but suffers from significantly lower accuracy. Interestingly, the accuracy of FedDF is lower than that of FedAvg, regardless of the model size. This result indicates that the soft targets generated by the logit ensemble are of insufficient quality to make meaningful progress in KD. FedDF can leverage KD without incurring extra communication cost. However, it runs KD on the server-side with inaccurate soft targets, which could ultimately harm model accuracy. DS-FL, the client-side KD-based FL method, adopts the logit ensemble and entirely avoids the model aggregation cost. Thus, the communication cost is only the number of classes $C$, but the model accuracy is significantly reduced.
@@
Computational Cost Analysis – We also compare the computational cost of our method to that of the same logit ensemble-based KD methods discussed in Table 2. Since FedDF and FedAvg are homogeneous FL methods that consider only a single type of model, a fair comparison of their computational costs with that of our method is not possible. Therefore, we focus on the training cost of $w_{l}$ only. Our method has the same computational cost as FedAvg (large) and FedDF (large), while achieving significantly higher accuracy. DS-FL is a heterogeneous FL method and its computational cost is directly comparable to our method. However, its accuracy is significantly lower than all the other methods due to the ineffective knowledge transfer through logit ensemble. This computational cost comparison suggests that our auxiliary model improves the target model’s accuracy without significantly increasing the total computational cost.
Method Comm. Cost Message Size Acc ($\%$)
FedAvg(small) $|w_{s}|$ 19K $57.47\pm 0.8\%$
@@ 3.4 Discussion on Impact of On-Device Knowledge Distillation on Generalization Performance
Here, we theoretically analyze the benefit of our design choices in terms of the target model’s generalization bound. Our analysis builds upon that of [8], extending it to account for the combination of labeled and unlabeled local data.
Problem Definition – We first define hypothesis $h:\mathcal{X}\rightarrow\mathcal{Y}$, with input space $\mathbf{x}\in\mathcal{X}$ and label space $y\in\mathcal{Y}$, and the hypothesis space $h\in\mathcal{H}$. The general loss function $l(h(\mathbf{x}),y)$ quantifies the classification error of the given hypothesis. The true global data distribution is defined as $\mathcal{D}$. Then, the true local data distribution from each client’s perspective is defined as $\mathcal{D}_{i},i\in[n]$. The data distribution of the limited number of local observations is defined as $\mathcal{\hat{D}}_{i}$. We also define the expected loss over an arbitrary data distribution $\mathcal{D}^{\prime}$ as $\mathcal{L}_{\mathcal{D^{\prime}}}(h)=\mathbb{E}_{\mathbf{x},y\sim\mathcal{D}^{\prime}}\left[l(h(\mathbf{x}),y)\right],\forall h\in\mathcal{H}$. Finally, we assume that $\mathcal{L}(h)$ is convex. In these settings, the generalization bound of the target model trained by our proposed method satisfies the following condition.
Now, to make our analysis self-contained, we state the two lemmas proposed in the previous work, which will be used to analyze the generalization of our proposed method.
Lemma 1 (Domain Adaptation [2]) With two true distributions $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, $\forall p\in(0,1)$ and hypothesis $\forall h\in\mathcal{H}$, with probability at least 1p1-p over the choice of samples, it satisfies
$\mathcal{L}_{\mathcal{D}_{1}}(h)\leq\mathcal{L}_{\mathcal{D}_{2}}(h)+\frac{1}{2}d(\mathcal{D}_{1},\mathcal{D}_{2})+v_{1},$ (3)
where $d(\mathcal{D}_{1},\mathcal{D}_{2})$ quantifies the distribution discrepancy between the two data distributions, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, $v_{1}$ is the minimal loss such that $v_{1}=\textrm{inf}_{h}\mathcal{L}_{\mathcal{D}_{1}}(h)+\mathcal{L}_{\mathcal{D}_{2}}(h)$.
@@
The first lemma measures the generalization bound for one data distribution based on the maximum discrepancy between the given two distributions. The second lemma elaborates the bound taking the limited training samples into account based on Hoeffding’s inequality. Our proposed method trains the target model $\mathbf{x}_{l}$ using $s<n$ strong clients only. Thus, we will analyze the generalization bound with respect only to the strong clients’ labeled data using the above two lemmas. Then, we will revise this bound concerning the on-device knowledge distillation with unlabeled local data.
Proposition 1 Consider $n$ clients in total and $s$ of them are strong clients. Each client has its own data distribution $\mathcal{D}_{i},\forall i\in[n]$. We define a combination of the labeled local data and the unlabeled local data as $\tilde{\mathcal{D}}_{i}$. For an arbitrary data distribution $\mathcal{D}^{\prime}$, $h_{\mathcal{D}^{\prime}}$ is the model best-trained on $\mathcal{D}^{\prime}$. Then, with probability of 1p1-p, our proposed on-device knowledge distillation-based FL guarantees the following generalization bound for the target model.
[…]
@@
Proof.
First, let us bound the loss over the true global data distribution $\mathcal{L}_{D}(\frac{1}{s}\sum_{i=1}^{s}h_{\mathcal{\hat{D}}_{i}})$ based on lemma 1. Note that the hypothesis is the model averaged only across $s$ strong clients.
$\mathcal{L}_{\mathcal{D}}\left(\frac{1}{s}\sum_{i=1}^{s}h_{\mathcal{\hat{D}}_{i}}\right)$ $\leq\frac{1}{s}\sum_{i=1}^{s}\mathcal{L}_{\mathcal{D}}\left(h_{\mathcal{\hat{D}}_{i}}\right)$
$\leq\frac{1}{s}\sum_{i=1}^{s}\left(\mathcal{L}_{\mathcal{D}_{i}}(h_{\mathcal{\hat{D}}_{i}})+\frac{1}{2}d(\mathcal{D}_{i},\mathcal{D})+v_{i}\right).$ (5)
Then, (5) can be further bounded using lemma 2 as follows.
[…]
Finally, the on-device knowledge distillation utilizes not only the strong clients’ labeled data but also the unlabeled data. Thus, we can simply replace the local dataset $\mathcal{\hat{D}}_{i}$ with the combination of the labeled and unlabeled data $\mathcal{\tilde{D}}_{i}$.
[…]
$\hskip 14.22636pt+\frac{1}{2s}\sum_{i=1}^{s}d(\mathcal{D}_{i},\mathcal{D})+\frac{1}{s}\sum_{i=1}^{s}v_{i}.$ (6)
Remark 1. This result highlights the key advantage of our method: on-device KD reduces the second term on the right-hand side of (6), leading to better generalization. Note that $|\hat{\mathcal{D}}_{i}|\leq|\tilde{\mathcal{D}}_{i}|$, where $\hat{\mathcal{D}}_{i}$ denotes the labeled data and $\tilde{\mathcal{D}}_{i}$ denotes the union of labeled and unlabeled data. Therefore, by expanding the data distribution, the proposed method can achieve a lower empirical loss.
Remark 2. As each client collects more unlabeled data, the size of $\mathcal{\tilde{D}}_{i}$ may increase and the bound decreases even further. This is a critical benefit in realistic FL environments because the new unlabeled data is most likely collected at the edge rather than the server. Thus, our method can be considered as a more scalable FL method especially when the strong clients can collect data on-the-fly. Note that in the server-side KD, the $\mathcal{\hat{D}}_{i}$ can only be the unlabeled data which is still smaller than $\mathcal{\tilde{D}}_{i}$.